My problem is to find the total surface area of a rectangular prism. The prisms dimensions are 4x by 3x for the base, and the prism is 5x units high. There is also a square hole in the prism that is in the middle, whose dimensions are x by 2x, and it is also 5x units high. Please help me!
Answer by mananth(16946)

The surface area of of rectangular prism is given by the formula
A= 2LW +2WH+2LH
Length = 4x
width = 3x
height = 5x
A= 2*3x*4x + 2*3x*5x + 2*4x*5x
A= 24x^2 +30x^2+40x^2
A= 84x^2
..
The area of the rectangular hole on the top surface and the base have to be subtracted.
area of rectangular hole on top and bottom =2*(2x * x)
=4x^2
..
The area on all four sides of the hollow prism have to be added.
= 2*2x*5x + 2*x*5x
=20x^2 +10x^2
=30x^2
total surface area = 84x^2-4x^2+30x^2
=110x^2



on